Prove that there is a positive integer that equals the sum of the positive integersnot exceeding it.
1
Expert's answer
2020-09-28T18:56:46-0400

Statement that need to be proved: there is a positive integer that equals the sum of the positive integers not exceeding it. In quantifiers this can be written as

"\displaystyle \exists \, n \in N\; (n = \sum_{i=0}^n i )"

We know that for finite sum "\displaystyle \sum_{i=0}^n i = \frac{n(n+1)}{2}"

So, "\displaystyle n = \frac{n(n+1)}{2}"

"n^2+n =2n"

"n^2-n=n(n-1)=0"

This has two solutions: "n=0" and "n=1". "n=0" doesn't satisfy the condition of being positive integer (by definition positive is >0, 0 is not greater than 0). Therefore, only "n=1" makes the statement true. We have showed that indeed there is a positive integer that equals the sum of the positive integers not exceeding it, and this positive integer is "n=1."